OpenStudy (honeymonster):

please help i'm really stuck :/ solve the equation..i'll post it in replies

OpenStudy (honeymonster):

\[3+ \frac{ 2-x }{ 2 }=\frac{ 2x+1 }{ 3 }\]

OpenStudy (precal):

get rid of the fractions and multiply it all by 6

OpenStudy (precal):